OpenStudy (ashking1):

Which equation is a quadratic equation? Answer y = 2x(5 - x) + 7x y = 8x + 2 y + x2 = (x - 2)(x + 8) y - 7x = (x2 + 1)(x - 3)

OpenStudy (lgbasallote):

simplify the choices A) \(y = 10x - 2x^2 + 7x\) B) \(y = 8x + 2\) C) \(y + x^2 = x^2 + 6x - 16\) D) \(y - 7x = x^3 - 3x^2 + x - 3\)

OpenStudy (lgbasallote):

which one is quadratic then?

OpenStudy (ashking1):

Is it A

OpenStudy (anonymous):

the equation having second degree

OpenStudy (ashking1):

is it A

OpenStudy (lgbasallote):

yep it's A

OpenStudy (anonymous):

yha! it is A

OpenStudy (lgbasallote):

C is possible BUT x^2 gets canceled so A is the only answer ^_^

OpenStudy (ashking1):

ok thks